OpenStudy (jpsmarinho):

\[ \lim_{x \to 0} \frac{5^x-1}{x} \] if I use the L'hospital rule, how I do? is this correct? \[ \lim_{x \to 0} \frac{ln(5^x-1)=lny}{lnx=lnz} \]

OpenStudy (anonymous):

the derivative of 5^x is \[ 5^x \ln 5\]

OpenStudy (anonymous):

\[ \lim_{x \to 0} \frac{5^x \ln 5}{1}= \ln 5\]

OpenStudy (jpsmarinho):

It's true @Spacelimbus, I'm wrong. Thanks! If, rather than \[5^x \] was \[x^x\]? I should use logarithm derivation right?

OpenStudy (anonymous):

yes but then things get pretty pretty weird, because the derivate of x^x is \[ x^x (\ln x +1)\] So you see already, for this expression, if x approaches zero, then this entire term isn't defined.

OpenStudy (anonymous):

and then you are stuck with l'Hospitals rule, because the x^x will never decay if you take it's derivative

OpenStudy (jpsmarinho):

I understand ^^. Thanks!